OpenStudy (anonymous):

Find the root:

OpenStudy (anonymous):

\[\sqrt{9-9i}\]

OpenStudy (anonymous):

solve z^2 = 9-9i

OpenStudy (anonymous):

@jayzdd How do you mean?

OpenStudy (anonymous):

you can use the trig form of the complex number z^2 = sqrt(9^2 + 9^2) ( cos(-45) + i sin(-45) )

ganeshie8 (ganeshie8):

you must be knowing that angles add up when you multiply two complex numbers

ganeshie8 (ganeshie8):

\[(r_1\angle \theta_1 )(r_2\angle \theta_2) = r_1r_2 \angle (\theta_1+\theta_2)\]

ganeshie8 (ganeshie8):

Finding \(\sqrt{9-9i}\) is same as finding a number which when multiplied by itself gives you \(9-9i\)

ganeshie8 (ganeshie8):

that is, finding some \(z\) such that \[z*z = 9-9i\]

ganeshie8 (ganeshie8):